# Robust Design and Evaluation of

### v5

@@ C.3 Estimation under proxy outcome bounds
Our next result characterizes the sharp bounds on the confounding function $\delta(x)$ under this assumption.
Lemma C.4.
@@
for $\underline{\delta}(x;\eta)=|1-\gamma_{1}(x)-\tilde{\mu}_{0}(x)|-\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=1-|\gamma_{1}-\tilde{\mu}_{0}|-\mu_{1}(x)$.
@@
Notice that the bounding functions stated in the main text immediately follow from Lemma C.4 if $\tilde{\mu}_{0}(x)\leq\gamma_{1}(x)$ and $\tilde{\mu}_{0}(x)+\gamma_{1}(x)\leq 1$ for all $x\in\mathcal{X}$. These restrictions are satisfied by the proxy outcomes used in BlattnerNelson(21) (see their Table 3) and MullainathanObermeyer(21) (see their Table 2) on average over the covariates.
We next extend our estimators for the conditional probability, positive class performance, and overall performance to bounding functions given in the main text based on a proxy outcome $\tilde{Y}_{i}$ that satisfies $\tilde{\mu}_{0}(x)\leq\gamma_{1}(x)$ and $\tilde{\mu}_{0}(x)+\gamma_{1}(x)\leq 1$ for all $x\in\mathcal{X}$:
@@
To extend our estimators, we observe that
@@
are uncentered efficient influence functions for $\mathbb{E}[\tilde{\mu}_{0}(X_{i})],\mathbb{E}[\gamma_{1}(X_{i})]$, $\mathbb{E}[\pi_{i}(X_{i})\tilde{\mu}_{0}(X_{i})]$, and $\mathbb{E}[\pi_{0}(X_{i})\gamma_{1}(X_{i})]$ respectively by standard calculations (e.g., Kennedy(22)-IFreview; Hines_2022).
To extend our estimators for the more general bounding functions given in Lemma C.4, we note that the lower bounding function can be written as $\min\{1-\gamma_{1}(x)+\tilde{\mu}_{0}(x),1-\gamma_{1}(x)+\tilde{\mu}_{0}(x)\}$ and the upper bounding function can be written as $\{1-\gamma_{1}(x)-\tilde{\mu}_{0}(x),\tilde{\mu}_{0}(x)+\gamma_{1}(x)-1\}$. We therefore can apply the same smooth approximations to the pointwise maximum and minimum functions that we discuss below in Appendix C.4.
C.3.1 Bounds on the conditional likelihood
We extend our nonparametric regression estimator for the proxy outcome bounds on the conditional probability. The construction is identical to Section 3, except we modify the pseudo-outcomes that are constructed in the second-stage. We again illustrate the procedure using two folds for simplicity. Using the first fold, we construct estimates $\hat{\eta}=\left(\hat{\pi}_{1}(\cdot),\hat{\mu}_{1}(\cdot),\widehat{\tilde{\mu}}_{0}(\cdot),\widehat{\gamma}_{1}(\cdot)\right)$. On the second fold, we construct the pseudo-outcome $\phi_{\mu,i}(\hat{\eta})+(1-D_{i})-\phi_{\pi\gamma,i}(\hat{\eta})-\phi_{\pi\tilde{\mu},i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})$ for the lower bound and the pseudo-outcome $\phi_{\mu,i}(\hat{\eta})+(1-D_{i})-\phi_{\pi\gamma,i}(\hat{\eta})+\phi_{\pi\tilde{\mu},i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})$ for the upper bound. We then regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified non-parametric regression procedure satisfying the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1).
Proposition C.1.
Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator. Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), and $\mathbb{P}(\epsilon\leq\hat{\pi}_{1}(X_{i})\leq 1-\epsilon)=1$ for some $\epsilon>0$. Define $\tilde{R}_{1}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, $\tilde{R}_{2}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\hat{\tilde{\mu}}_{0}(X_{i})-\tilde{\mu}_{0}(X_{i}))\mid X_{i}=x]$, $\tilde{R}_{3}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\hat{\gamma}_{1}(X_{i})-\gamma_{1}(X_{i}))\mid X_{i}=x]$, and $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$. Then,
@@ C.3.1 Bounds on the conditional likelihood
C.3.2 Bounds on positive class performance
We analogously estimate the proxy outcome bounds on positive class performance by solving sample linear-fractional programs as in Section 4.1 of the main text. We again split the data into three folds. On the first fold, we construct nuisance function estimators $\hat{\eta}$. On the second fold, we construct the pseudo-outcomes $\phi_{\mu,i}(\hat{\eta})+(1-D_{i})-\phi_{\pi\gamma,i}(\hat{\eta})-\phi_{\pi\tilde{\mu},i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})$ for the lower bound and the pseudo-outcome $\phi_{\mu,i}(\hat{\eta})+(1-D_{i})-\phi_{\pi\gamma,i}(\hat{\eta})+\phi_{\pi\tilde{\mu},i}(\hat{\eta})-\phi_{\pi\mu,i}(\hat{\eta})$ for the upper bound. We then regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified non-parametric regression procedure, yielding estimators $\widehat{\underline{\delta}}(X_{i})$, $\widehat{\overline{\delta}}(X_{i})$ respectively. Finally, on the third fold, we estimate the upper bound on positive class performance by solving
where now $\widehat{\Delta}_{n}=\{(1-D_{i})\widehat{\underline{\delta}}(X_{i})\leq\tilde{\delta}_{i}\leq(1-D_{i})\widehat{\overline{\delta}}(X_{i})\}$. Our estimator for the lower bound is defined analogously. These estimators are again equivalent to solving linear programs by the Charnes-Cooper transformation.
Again, using the same arguments as in the proof of Proposition 4.1 we can show that these estimators converge quickly to the proxy outcome bounds on positive class performance. We next state the result, but skip the proof for brevity.
Proposition C.2.
Define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$, $R_{2,n}=\|\tilde{\hat{\mu}}_{0}(\cdot)-\tilde{\mu}_{0}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$, and $R_{3,n}=\|\hat{\gamma}_{1}(\cdot)-\gamma_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$, (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1}(X_{i})\geq\epsilon)=1$, (iv) there exists $c>0$ such that $\mathbb{P}(\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})])=1$; (v) $\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|=o_{\mathbb{P}}(1),\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{\mathbb{P}}(1),\|\tilde{\hat{\mu}}_{0}(\cdot)-\tilde{\mu}_{0}(\cdot)\|=o_{\mathbb{P}}(1)$, and $\|\hat{\gamma}_{1}(\cdot)-\gamma_{1}(\cdot)\|=o_{\mathbb{P}}(1)$; and (vi) $\widehat{\Delta}_{n}$ is non-empty with probability one. Then,
$\left|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)-\overline{\mbox{perf}}_{+}(s;\beta)\right|=O_{\mathbb{P}}\left(\frac{1}{\sqrt{n}}+(R_{1,n}+R_{2,n}+R_{3,n})+\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}+\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}\right),$
and the analogous result holds for $\widehat{\underline{\mbox{perf}}}_{+}(s;\beta)$.
Of course, if the nonparametric regression procedure further satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), we can further control the errors $\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}$, $\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}$ as in Corollary 4.1 of the main text.
C.3.3 Bounds on overall performance
@@
The lower bound can be written analogously. Like in our analysis of nonparametric outcome bounds in Section 4.2, the bounds on overall performance are again the expectation of some known function of the data and nuisance parameters $\eta=(\mu_{1}(\cdot),\pi_{0}(\cdot),\tilde{\mu}_{0}(\cdot),\gamma_{1}(\cdot))$. We can therefore construct debiased estimators as in the main text.
In particular, we again randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\hat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
@@
Proposition C.3.
@@
C.4 Estimation under instrumental variable bounds
We now extend our estimators for the bounds on overall performance, positive class performance and the conditional probability to bounding functions based on an instrumental variable $Z_{i}$ with finite support, as discussed in Section 2.3.3.
As notation, let $p=|\mathcal{Z}|$ denote the number of unique instrument values, $\lambda_{z}(x):=\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]$ and $\kappa_{z}(x):=\mathbb{P}(D_{i}=0\mid X_{i}=x,Z_{i}=z)$. Recall that for any value of the instrument $z\in\mathcal{Z}$, the confounding function satisfies, for all $x\in\mathcal{X}$,
@@ C.4 Estimation under instrumental variable bounds
are the uncentered efficient influence functions for $\mathbb{E}[\lambda_{z}(X_{i})]$ and $\mathbb{E}[\kappa_{z}(X_{i})]$ respectively.
C.4.1 Bounds on the conditional likelihood
Finally, we briefly extend our nonparametric regression estimator for the instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$. We again illustrate the procedure using two folds for simplicity. Using the first fold, we construct nuisance function estimates $\hat{\eta}=(\hat{\lambda}_{z}(\cdot),\hat{\kappa}_{z}(\cdot),\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot))$. On the second fold, we construct the pseudo-outcome $\phi_{\kappa,z,i}(\hat{\eta}),\phi_{\lambda,z,i}(\hat{\eta})$ for the upper bound and $\phi_{\lambda,z,i}(\hat{\eta})$ for the lower bound. We regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure satisfying the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1).
@@ C.4.1 Bounds on the conditional likelihood
Proposition C.4.
Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1) and $\mathbb{P}(\epsilon\leq\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))=1$ for some $\epsilon>0$. Define $\tilde{R}_{2}(x)=\widehat{\mathbb{E}}_{n}[(\mathbb{P}(Z_{i}=z\mid X_{i}=x)-\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))(\hat{\lambda}_{z}(x)-\lambda_{z}(x))\mid X_{i}=x]$, $\tilde{R}_{3}(x)=\widehat{\mathbb{E}}_{n}[(\mathbb{P}(Z_{i}=z\mid X_{i}=x)-\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))(\hat{\kappa}_{z}(x)-\kappa_{z}(x))\mid X_{i}=x]$, and $R^{2}_{oracle}(z)=\mathbb{E}[\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$. Then,
@@
C.4.2 Bounds on positive class performance
Under instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$, we estimate the upper bound on positive class performance by solving sample linear-fractional programs for an alternative constraint set. We again split the data into three folds. Using the first fold, we construct nuisance function estimators $\hat{\eta}_{-k}=(\hat{\pi}_{1,-k}(\cdot),\hat{\mu}_{1,-k}(\cdot),\hat{\lambda}_{z,-k}(\cdot),\hat{\kappa}_{z,-k}(\cdot),\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot))$. Using the second fold, we construct the pseudo-outcome $\phi_{\kappa,z,i}(\hat{\eta}),\phi_{\lambda,z,i}(\hat{\eta})$ for the upper bound and $\phi_{\lambda,z,i}(\hat{\eta})$ for the lower bound. We regress the estimated pseudo-outcomes on the covariates $X_{i}$ using a researcher-specified nonparametric regression procedure, yielding estimators $\widehat{\overline{\delta}}(X_{i}),\widehat{\underline{\delta}}(X_{i})$ respectively. Finally, on the third fold, we construct the fold-specific estimate of the upper bound on positive class performance by solving
where now $\widehat{\Delta}_{n}=\{(1-D_{i})\widehat{\underline{\delta}}(X_{i})\leq\tilde{\delta}_{i}\leq(1-D_{i})\widehat{\overline{\delta}}(X_{i})\}$. Our estimator for the lower bound is defined analogously. These estimators are again equivalent to solving $K$ linear programs by the Charnes-Cooper transformation.
Again, using the same arguments as in the proof of Proposition 4.1 we can show that these estimators converge quickly to the proxy outcome bounds on positive class performance. We next state the result, but skip the proof for brevity.
Proposition C.5.
Define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$, $R_{2,n}=\|\hat{\lambda}_{z}(\cdot)-\lambda_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$ and $R_{3,n}=\|\hat{\kappa}_{z}(\cdot)-\kappa_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}\{\mathbb{P}(Z_{i}=z\mid X_{i})\geq\delta\}=1$ and $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$; (iii) there exists $\epsilon>0$ such that {^$\mathbb{P}\{\hat{\mathbb{P}}(Z_{i}=z\mid X_{i})\geq\epsilon\}=1$ and $\mathbb{P}(\hat{\pi}_{1}(X_{i})\geq\epsilon)=1$; (iv) there exists some $c>0$ such that $\mathbb{P}(\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta}_{-k})+\widehat{\underline{\delta}}_{i}]>c)=1$; (v) $\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\pi}_{1}-\pi_{1}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\lambda}_{z}(\cdot)-\lambda_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\kappa}_{z}(\cdot)-\kappa_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, and ^$\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i})-\mathbb{P}(Z_{i}=z\mid X_{i})\|=o_{\mathbb{P}}(1)$; and (vi) $\widehat{\Delta}_{n}$ is non-empty with probability one. Then,
$\left|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)-\overline{\mbox{perf}}_{+}(s;\beta)\right|=O_{\mathbb{P}}\left(1/\sqrt{n}+(R_{1,n}+R_{2,n}+R_{3,n})+\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}+\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}\right),$
and the analogous result holds for $\widehat{\underline{\mbox{perf}}}_{+}(s;\beta)$.
Of course, if the nonparametric regression procedure further satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption B.1), we can further control the errors $\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}$, $\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}$ as in Corollary 4.1 of the main text.
C.4.3 Bounds on overall performance
Under instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$, the upper bound on overall performance can be written as

### v6

@@ D.1 Estimation under Proxy Outcome Bounds
Equation 14 states that the probability the true and proxy outcomes agree is the same regardless of selection status, conditional on covariates. In consumer lending, this says that default on credit cards is equally predictive of default on personal loans for both funded and unfunded applicants with the same credit profile.
Our next result characterizes the sharp bounds on the confounding function $\delta(x)$ under Equation 14.
Lemma D.1.
@@
for $\underline{\delta}(x;\eta)=|1-\gamma_{1}(x)-\tilde{\mu}_{0}(x)|-\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=1-|\gamma_{1}(x)-\tilde{\mu}_{0}(x)|-\mu_{1}(x)$.
@@
In many applications, the proxy satisfies additional restrictions that simplify the bounding functions. The bounding functions stated in Equation 6 of the main text immediately follow from Lemma D.1 if further $\tilde{\mu}_{0}(x)\leq\gamma_{1}(x)$ and $\tilde{\mu}_{0}(x)+\gamma_{1}(x)\leq 1$ for all $x\in\mathcal{X}$. These restrictions are satisfied by the proxy outcomes used in BlattnerNelson(21) (see their Table 3) and MullainathanObermeyer(21) (see their Table 2) on average over the covariates.
For the remainder of this section, we extend our estimators for the conditional probability, positive class performance, and overall performance to the simplified bounding functions based on a proxy outcome $\tilde{Y}_{i}$ satisfying $\tilde{\mu}_{0}(x)\leq\gamma_{1}(x)$ and $\tilde{\mu}_{0}(x)+\gamma_{1}(x)\leq 1$ for all $x\in\mathcal{X}$. The simplified bounding functions are:
@@
We observe that
@@
are influence functions for $\mathbb{E}[\tilde{\mu}_{0}(X_{i})],\mathbb{E}[\gamma_{1}(X_{i})]$, $\mathbb{E}[\pi_{i}(X_{i})\tilde{\mu}_{0}(X_{i})]$, and $\mathbb{E}[\pi_{0}(X_{i})\gamma_{1}(X_{i})]$ respectively by standard calculations (e.g., Kennedy(22)-IFreview; Hines_2022).
To extend our estimators to the general proxy outcome bounds in Lemma D.1, we use smooth approximations to the absolute value functions. The lower bounding function can also be written as $\max\{1-\gamma_{1}(x)-\tilde{\mu}_{0}(x),\tilde{\mu}_{0}(x)+\gamma_{1}(x)-1\}-\mu_{1}(x)$ and the upper bounding function as $\min\{1-\gamma_{1}(x)+\tilde{\mu}_{0}(x),1+\gamma_{1}(x)-\tilde{\mu}_{0}(x)\}$. We apply the log-sum-exponential function to approximate these pointwise maximum and minimum operations as we discuss below in Appendix D.2. Alternatively, for the overall performance estimator, we could invoke a margin condition as discussed in LevisEtAl(23)-CovariateAssistedIV; semenova2023adaptive.
D.1.1 Bounds on the Conditional Likelihood
We extend our nonparametric regression estimator for the proxy outcome bounds on the conditional probability. The construction is identical to Section 3, except we modify the pseudo-outcomes that are constructed in the second-stage. We again illustrate the procedure using two folds for simplicity.
Estimation Procedure for Proxy Outcome Bounds:
[…]
Bound on Integrated Mean Square Error Convergence:
For modified definitions of the smoothed doubly robust residuals and smoothed bias, we again derive a bound on the integrated mean square error of our feasible estimators relative to that of an infeasible oracle nonparametric regression that has access to the true nuisance functions using the same arguments as Proposition 3.1. We state the result for completeness, but skip the proof for brevity.
Proposition D.1.
Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator. Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption A.1), and $\mathbb{P}(\epsilon\leq\hat{\pi}_{1}(X_{i})\leq 1-\epsilon)=1$ for some $\epsilon>0$. Define $\tilde{R}_{1}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, $\tilde{R}_{2}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\hat{\tilde{\mu}}_{0}(X_{i})-\tilde{\mu}_{0}(X_{i}))\mid X_{i}=x]$, $\tilde{R}_{3}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\hat{\gamma}_{1}(X_{i})-\gamma_{1}(X_{i}))\mid X_{i}=x]$, and $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$. Then,
@@ D.1.1 Bounds on the Conditional Likelihood
D.1.2 Bounds on Overall Performance
@@
The lower bound can be written analogously. Like in Section 4.1 of the main text, both bounds are linear functionals of known functions of the data and identified nuisance parameters. This linearity enables us to construct debiased estimators using standard arguments.
Estimation Procedure for Proxy Outcome Bounds:
We sketch the construction of our estimators based on $K$-fold cross-fitting. We randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\hat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
@@
Consistency and Asymptotic Normality:
Using the same arguments as in the proof of Proposition 4.1, we can again show that these estimators converge to the proxy outcome bounds on overall performance and are jointly asymptotically normal. We next state the result, but skip the proof for brevity.
Proposition D.2.
@@
$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,\Sigma\right)$
D.1.3 Bounds on Positive Class Performance
We analogously estimate the proxy outcome bounds on positive class performance by solving sample linear fractional programs as in Section 4.2 of the main text. Building on Proposition 4.2, all that needs to be modified is the choice estimator for the bounding functions. As an illustration, we sketch out how the sample linear fractional program can be combined with our pseudo-regression procedure for proxy outcome bounds discussed earlier in Appendix D.1.1.
Estimation Procedure for Proxy Outcome Bounds:
@@ D.1.3 Bounds on Positive Class Performance
where now $\widehat{\Delta}_{n}=\{(1-D_{i})\widehat{\underline{\delta}}(X_{i})\leq\tilde{\delta}_{i}\leq(1-D_{i})\widehat{\overline{\delta}}(X_{i})\}$. These estimators are again equivalent to solving linear programs by the Charnes-Cooper transformation. Proposition 4.2 can be applied, and this estimator has the same partial double robustness property. The error of this estimator again depends on the root mean square error of the estimated bounding functions. If the nonparametric regression procedure further satisfies the $L_{2}(\mathbb{P})$-stability condition, then we can further control these errors using the results in Appendix D.1.1.
D.2 Estimation under Instrumental Variable Bounds
Suppose we observe an instrumental variable $Z_{i}\in\mathcal{Z}$ with finite support that generates quasi-random variation in selection decisions but does not directly affect outcomes as discussed in Section 2.3.3. The instrument provides identifying power under two assumptions: (i) $Z_{i}$ exogenously shifts the selection decision but not the true outcome, $Y_{i}^{*}\perp\!\!\!\perp Z_{i}\mid X_{i}$; and (ii) $Z_{i}$ is relevant, meaning there exists $z,z^{\prime}\in\mathcal{Z}$ with $\mathbb{P}(D_{i}=1\mid X_{i}=x,Z_{i}=z)\neq\mathbb{P}(D_{i}=1\mid X_{i}=x,Z_{i}=z^{\prime})$. Under these conditions, classic results by Manski(94)-SampleSelectionBounds imply bounds on the confounding function. As notation, let $p=|\mathcal{Z}|$ denote the number of unique instrument values, $\lambda_{z}(x):=\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]$ and $\kappa_{z}(x):=\mathbb{P}(D_{i}=0\mid X_{i}=x,Z_{i}=z)$.
Recall that for any value of the instrument $z\in\mathcal{Z}$, the confounding function satisfies, for all $x\in\mathcal{X}$,
@@ D.2 Estimation under Instrumental Variable Bounds
are the influence functions for $\mathbb{E}[\lambda_{z}(X_{i})]$ and $\mathbb{E}[\kappa_{z}(X_{i})]$ respectively.
D.2.1 Bounds on the Conditional Likelihood
We briefly extend our nonparametric regression estimator for for the instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$. The construction is identical to Section 3, except we modify the pseudo-outcomes that are constructed in the second-stage. We again illustrate the procedure using two folds for simplicity.
Estimation Procedure for Single-Instrument Bounds:
[…]
Bound on Integrated Mean Square Error Convergence:
@@ D.2.1 Bounds on the Conditional Likelihood
Proposition D.3.
Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption A.1) and $\mathbb{P}(\epsilon\leq\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))=1$ for some $\epsilon>0$. Define $\tilde{R}_{2}(x)=\widehat{\mathbb{E}}_{n}[(\mathbb{P}(Z_{i}=z\mid X_{i}=x)-\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))(\hat{\lambda}_{z}(x)-\lambda_{z}(x))\mid X_{i}=x]$, $\tilde{R}_{3}(x)=\widehat{\mathbb{E}}_{n}[(\mathbb{P}(Z_{i}=z\mid X_{i}=x)-\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))(\hat{\kappa}_{z}(x)-\kappa_{z}(x))\mid X_{i}=x]$, and $R^{2}_{oracle}(z)=\mathbb{E}[\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$. Then,
@@
D.2.2 Bounds on Overall Performance
Under instrumental variable bounds at a single instrument value $z\in\mathcal{Z}$, the upper bound on overall performance can be written as